Show that the values are equal.
1-2 sin^2 30 degrees= cos 60 degrees

Answer by eperette(173)

1-2sin^2(30)=cos(60)
1-2sin^2(30)= 1-2(sin(30))^2 = 1-2(1/2)^2=1-2(1/4)=1-2/4=2/4=1/2 = cos(60)